OpenStudy (destinyjohnson):

Can someone help? Michelle draws a card from a standard deck of 52 cards. She replaces the card and draws a second card. What is the probability that she draws a heart and then a spade? I think it's either 13/52 or 26/52

OpenStudy (jango_in_dtown):

Yeah I can help

OpenStudy (destinyjohnson):

Thank you.

OpenStudy (jango_in_dtown):

So how many hearts are there in a deck of cards?

OpenStudy (destinyjohnson):

13

OpenStudy (jango_in_dtown):

And how many spade?

OpenStudy (destinyjohnson):

The same, 13.

OpenStudy (jango_in_dtown):

So let A denote the event that a heart is drawn and B denote the event that a spade is drawn. Then P(A)=13/52=1/4 and P(B)=13/52=1/4. Required probability=P(AB)=P(A)P(B) since the events are independent =(1/4)*(1/4)=1/16

OpenStudy (jango_in_dtown):

@DestinyJohnson

OpenStudy (destinyjohnson):

so the answer is 1/16?

OpenStudy (jango_in_dtown):

yes

OpenStudy (destinyjohnson):

Wow, I was way off xD, thank you for your help though. I appreciate it.